OpenStudy (anonymous):

Help? :c Solve: x2 − 2x − 24 = 0. x = 4, x = 6 x = 4, x = −6 x = −4, x = 6 x = −4, x = −6

OpenStudy (anonymous):

x2 − 2x − 24=x2 − 6x +4x− 24=(x-6)(x+4)

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

well done @jeanelle417